XYZ made a number of sales in 2014 and 2015 that required use of the instalment method. The following information regarding the sales is available:

Instalment Sales 150000 160000
Cost Of Instalment Sales 115500 104000
Collections on 2014 Sales 40000 70000
Collections on 2015 Sales 30000


Give the journal entries relating to instalment sales for the years 2014 and 2015. Ignore interest charges. Please make sure your final answer(s) are accurate to 2 decimal places. Assume that all entries are made on December 31. Enter an appropriate description when entering the transactions in the journal. Dates must be entered in the format dd/mmm (ie. January 15 would be 15/Jan).
Answer #1

Journal entries for 2014 sales and its cash collections

Date Account Titles Debit Credit
2014
31-Dec Instalment Accounts Receivable 150000
       Inventory 115500
       Deferred Gross Profit 34500
(Entry for initial recognition of instalment sale)
31-Dec Cash 40000
      Instalment Accounts Receivable 40000
(Cash Received for sales)
Cost of Goods Sold 30800
Deferred Gross Profit 9200
        Sales Revenue 40000
(Recognition of sales revenue for part of cash collected)
2015
31-Dec Cash 70000
      Instalment Accounts Receivable 70000
(Cash Received for sales)
Cost of Goods Sold 53900
Deferred Gross Profit 16100
        Sales Revenue 70000
(Recognition of sales revenue for part of cash collected)

Cost of Sales Percentage = Cost of Sales / Instalment sales x 100
= 115500 / 150000 x 100 = 77%

Sales is recognised when cash is collected, cost of goods sold for instalment sales = Cash Collected x 77%
Deferred Gross Profit written back = Cash Collected - Cost of Goods Sold

Journal entries for 2015 sales and its cash collections

Date Account Titles Debit Credit
2015
31-Dec Instalment Accounts Receivable 160000
       Inventory 104000
       Deferred Gross Profit 56000
(Entry for initial recognition of instalment sale)
31-Dec Cash 30000
      Instalment Accounts Receivable 30000
(Cash Received for sales)
Cost of Goods Sold 19500
Deferred Gross Profit 10500
        Sales Revenue 30000
(Recognition of sales revenue for part of cash collected)

Cost of Sales Percentage = Cost of Sales / Instalment sales x 100
= 104000 / 160000 x 100 = 65%